OpenStudy (16shuston):

he school production of​ 'Our Town' was a big success. For opening​ night, 491 tickets were sold. Students paid ​$1.50 ​each, while​ non-students paid $3.50 each. If a total of $ 956.50 was​ collected, how many students and how many​ non-students attended solve the system x+y = 491 2.5x+4.5y= 956.50 x= ? y= ?

OpenStudy (anonymous):

What did you get for your answers?

OpenStudy (16shuston):

im not sure where to go from here

OpenStudy (anonymous):

Solve this system of equations. x+y = 491 2.5x+4.5y= 956.50

OpenStudy (16shuston):

how?

OpenStudy (anonymous):

There are a few ways. What way you want to use. Your teacher has probably showed you a few ways to solve system of equations. You have elimination substitution. The subsitution would be to solve for y in 2.5x+4.5y= 956.50, which would be 2.5x+4.5y= 956.50 2.5x+4.5y= 956.50- 2.5x 2.5x-2.5x+4.5y= 956.50- 2.5x 4.5y= 956.50- 2.5x \[\large \frac{4.5y}{4.5} = \frac{956.50-2.5}{4.5} \] \[\large y= \frac{956.50-2.5}{4.5} \] \[\large y= 212.55-.55 \] Now take Y and substitute it into the top equation x+y=491 x+(212.55-.55)=491 Now turn those decimals to whole numbers by multiplying by 100 on both sides of the equation 100(x+(212.55-.55))=491 * 100 100x+21255-55=49100 Now just solve for x from this equation 100x+21255-55=49100. Can you solve x now?

OpenStudy (16shuston):

yes thank you so much it makes more sense now.

OpenStudy (anonymous):

Once you find x just put x into the equation x+y=491 and solve for y

OpenStudy (16shuston):

okay